a piece of bone from an organism is found to contain 10% of the carbon -14 is 5730 years then how long ago was the organism was alive?

please help me solve!!!!!
i really dont know how tio solve pleaseeeee help.....

Answer by scott8148(6628)

the half-life of carbon-14 is 5730yr
this means that in 5730yr, 50% of the original amount will remain
in another 5730yr, 25% will remain
and so on

the number of half-lives can be found from ___ p = (1/2)^n
where p is the fraction remaining and n is the number of half-lives

taking the log ___ log(p) = n * log(1/2)
dividing by log(1/2) ___ [log(p)] / [log(1/2)] = n

[log(.10)] / [log(1/2)] = n
using a calculator ___ n = 3.32 (approx)

so the time is ___ 3.32 * 5730 = 19024yr